The perimeter of a rectangle is 72 m. If the width were doubled and the lengh were increased by 9 m, then the perimeter would be 114 m. What are the dimensions?
Answer by TheProdicalSon(34)

2w + 2l = 72
2h + 2l+18 = 114
-18 -18
2h + 2l = 96
(96 - 72 = 24)
2h = 24
h = 12
w = 12
24 + 2l = 72
2l = 48
l = 24

length is 24
width is 12